Question 2 2.1 Derive the following O(h) three-point formula (including the error term): f(x) = f(x, +h)-f(xo – h))f wherePLEASE ASSIST WITH BOTH OF THEM
Answer #1


2! 3! [. 12 And, ち 3! 2! solution Let us consider, Taylor polynomial tell three differentiation term f(xoth) = f(x0) + f(x0)(2) Given, #(x) = x enca) = f(x) = ax 1 + enca) f(X) 1 + 2 (1) ş Ex aet value of f(x) et 228,3 is (803) = 1 t en (803) f(. #(20-h) = $(813-015) f(x0-6) H708) = 708 en (708) = 16.0221 And f(20) = R. > {(8-3) = -0.0145 ::$(813) 19:1378 – 16.022